OpenStudy (anonymous):

match the given polynomial to the one written in quadratic form. can someone please explain how to do this??

OpenStudy (anonymous):

A quadratic is an equation in this format. \[Y = Ax ^{2}+Bx+C\] If the problem gives you a polynomial, it may look like \[Y = I(Jx-K)^{2}\] just carry out the squaring of the part in parenthesis and multiply everything by I to get the quadratic format.

OpenStudy (anonymous):

sooo y=I(Jx-k)^2 would be I*jx^2-k^2??

OpenStudy (anonymous):

@Matt.Mawson

OpenStudy (anonymous):

nope. Try expanding the square and doing each multiplication individualy (term 1 times term 3 and term 1 times term 4, then term 2 times term 3 and term 2 times term 4, etc.\[Y = I \times (Jx - K)(Jx-K)\]

OpenStudy (anonymous):

but they are all different variables.

OpenStudy (anonymous):

True, but try to think of I, J, and K as constants, since they represent the coefficients of our function. Each function has coefficients, but they are usually small whole numbers.

OpenStudy (anonymous):

okay so like x^4-8x^2+12 would be (x^2)^2-8x^2+12?

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

so if I wanted to figure out a solution like it is telling me to with separate problems, would I just enter in the given roots or is there a way to figure it out?

OpenStudy (anonymous):

@Matt.Mawson